OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1)(greater than or equal to) 4(2x - 3) - 3? (5 points) A) x (less than or equal to) 6 B) 6 (greater than or equal to) x C) A number line with a closed circle on 6 and shading to the left. D) A number line with a closed circle on 6 and shading to the right.

OpenStudy (anonymous):

3(2x-1)>=4(2x-3)-3 6x-3>=8x-12-3 6x-3>=8x-15 6x-8x>=-15+3 -2x>=-12 2x<=12 x<=6 or 6 (greater than or equal to) x

OpenStudy (anonymous):

Thanks(:

OpenStudy (anonymous):

not a way ......its d